OpenStudy (anonymous):

if M is the midpoint of line AB, find b. Then use distance to justify your answers. A(5,1), M(-3,7)

OpenStudy (anonymous):

use desmos to graph this

OpenStudy (anonymous):

desmos?

OpenStudy (anonymous):

graphing calculator

OpenStudy (ddcamp):

Desmos is a great online graphing tool, but its not a good idea to become too dependent on it.

OpenStudy (anonymous):

im not suppose to graph it

OpenStudy (ddcamp):

Let's start by finding the midpoint using the midpoint equation(s): \[x = \frac{x_1 + x_2}{2} \\ \\y = \frac{y_1+y_2}{2}\]

OpenStudy (anonymous):

i need to find point B using the distance formula or midpoint

OpenStudy (anonymous):

Ok !

OpenStudy (anonymous):

(1,4) is the midpoint

OpenStudy (anonymous):

and 10 is the distance

OpenStudy (anonymous):

yeah lol i got that already

OpenStudy (ddcamp):

OK. good!

OpenStudy (ddcamp):

For the part where you "use distance to justify your answer," find the distance between the two endpoints, then find the distance between an endpoint and the midpoint. If the second distance is half the first, then you probably got the midpoint right!

OpenStudy (anonymous):

okay but what is point B?

OpenStudy (anonymous):

yo @Robert136 come help us bruuh

OpenStudy (anonymous):

Think *brian*

OpenStudy (anonymous):

what grade u in @cali16

OpenStudy (anonymous):

just started 12th today

OpenStudy (anonymous):

DANG thats why i cant get this

OpenStudy (ddcamp):

*doh* I read the question backwards, sorry.

OpenStudy (ddcamp):

First, to find B, let's find the change in x coordinates between A and M.

OpenStudy (anonymous):

like the distance??

OpenStudy (ddcamp):

Yeah, like the distance but just in the x direction.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

dang u fast bruuh @DDCamp

OpenStudy (ddcamp):



OpenStudy (anonymous):

Yeah bruuh thats what i got so u got this right cuz i got this girl on my back man

OpenStudy (anonymous):

so do you add them up to get B?? for the midpoint i got (1,4)

OpenStudy (anonymous):

see yall GL @cali16 hope u understand it

OpenStudy (ddcamp):

So M is the midpoint between A and B, but we already know A and M. Finding the midpoint between A and M would be finding this point: